2. Find the general solution of the following non-homogeneous differential equation dy dy sin (2t) (2) 2 +y= dt dt2 Now, letFind the general solution of the following non-homogeneous differential equation d 2 y dt2 + 2 dy dt + y = sin (2t). (2) Now, let y(t) be the general solution you find, when happen if we take lim t→+∞ y(t)?
Answer #1

And differential Given equation non- homogeneous is 2 + 2 dy + y = sin (24) - (6² +2D+1) = sin (21) where Deut Auxiliary equa(p. 1.) = (2D+3) - (20+3)(20-3) Sin (24) (2D+3) (4D²-9) Sin (24) (2D+3) Sin (24) -4.2² -9 (2D+3) Sin (24) -16-9 (2D+3) Sin (2